A 0.42 kg soccer ball moving at 12 m/s is kicked with a force of 126 N for 0.02 seconds. (a) calculate the impulse on the ball. (b) calculate the change in momentum on your foot. (c) what is the final velocity ofthe ball


1
Expert's answer
2021-01-19T07:12:01-0500

a) We can find the impulse on the ball from the definition of impulse:


"J=F\Delta t=126\ N\cdot0.02\ s=2.52\ N\cdot s."

b) Since the impulse on the ball equals the change in momentum on foot, we can write:


"J=\Delta p_{foot}=2.52\ N\cdot s."

c) Let's first calculate the change in velocity of the ball from the impulse-momentum change equation:


"F\Delta p=m\Delta v,""\Delta v=\dfrac{F\Delta p}{m}=\dfrac{J}{m}=\dfrac{2.52\ N\cdot s}{0.42\ kg}=6\ \dfrac{m}{s}."

Finally, we can find the final velocity of the ball:


"\Delta v=v_f-v_i,""v_f=\Delta v+v_i=6\ \dfrac{m}{s}+12\ \dfrac{m}{s}=18\ \dfrac{m}{s}."

Answer:

a) "J=2.52\ N\cdot s."

b) "\Delta p_{foot}=2.52\ N\cdot s."

c) "v_f=18\ \dfrac{m}{s}."